OpenStudy (anonymous):

|15 8| - |2 5| = |13 3| |4x 0| |x 3x+6| |12 4y+8| Find the values of the variable

OpenStudy (ivettef365):

4x-x= 3x since 3x = 12 then x = 4

OpenStudy (ivettef365):

Therefore 3x+6 = 3(4) + 6 therefore it is 18 and 0 - 18 = -18

OpenStudy (anonymous):

isnt y=2 or -2?

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

I have 4y+8=-18 4y=-26 y= -6.5

OpenStudy (anonymous):

\[\left[\begin{matrix} 15 & 8 \\ 4x & 0\end{matrix}\right]-\left[\begin{matrix}2 & 5 \\ x &3x+6\end{matrix}\right]\]=\[\left[\begin{matrix}13 & 3 \\12 &4y+8\end{matrix}\right]\] \[\left[\begin{matrix}15-2 & 8-5 \\4x-x & 0-3x-6\end{matrix}\right]=\left[\begin{matrix}13 & 3 \\12 &4y+8\end{matrix}\right]\] \[\left[\begin{matrix}13 & 3 \\3x & -3x-6\end{matrix}\right]=\left[\begin{matrix}13 & 3 \\ 12 & 4y+8\end{matrix}\right]\] 3x=12 -3x-6=4y+8 Find the values of x &y